OpenStudy (nicholausblackmon):

Write an equation in slope-intercept form of the line through point P(6,7) with slope -2. Will medal and fan!

OpenStudy (nicholausblackmon):

@dan815

OpenStudy (nicholausblackmon):

i think it is y=-2x+19, what about you?

OpenStudy (nicholausblackmon):

@ganeshie8

OpenStudy (nicholausblackmon):

The options are y-7=-2(x-6) y=-2+7 Y=-2x+19 Y-6=-2(x-7)

OpenStudy (swamijatin2):

1st option is right

OpenStudy (nicholausblackmon):

how is it right?

OpenStudy (swamijatin2):

by righting the equation y-y1=m(x-x1)

OpenStudy (nicholausblackmon):

Ohhhh! i forgot about the formula!

OpenStudy (swamijatin2):

y-7=-2(x-6)

OpenStudy (swamijatin2):

now u got it